OpenStudy (anonymous):

Please help???????????????????? Describe why compound interest earns more money than simple interest

OpenStudy (anonymous):

simple interest: you invest 100 dollars at 10% interest for 3 years. after 3 years, your return is (10+10+10)=30 and total return: 100(original investment)+30=130. so, the future value of the 100$ investment at 10% simple interest for 3 years is 130$. FV (simple interest)= Original Investment x (1+(interest rate*number of years) =100* (1 +(0.1+3)) =130$ compound interest: you invest 100 dollars at 10% compound interest for 3 years. for first year, 10% of 100$ is 10$, for the second year the interest will be 10% plus additional interest on the 10% of the first year, and again 10% interest on the third year plus interest on the previous year's interest. so, compound interest basically means interest on interest. FV(Compound interest)= Original Investment x ((1+interest rate)^number of years). =100$ * (1 + 0.1)^3 =1331$